-4x^2=32x+ 48

answer : x=-6 or -2
need a break down of why the answer is the answer

Answer by checkley77(12844)

-4x^2=32x+48
-4x^2-32x-48=0
-4(x^2+8x+12)=0
-4(x+6)(x+2)=0
x+6=0
x=-6 ans.
x+2=0
x=-2 ans.